# Scene Graph Generation in Large-Size VHR Satellite Imagery: A Large-Scale Dataset and A Context-Aware Approach

### v1

@@ Abstract
Scene graph generation (SGG) benchmark, large-size satellite imagery, object detection, relationship prediction.
@@ 2 Related Work
In this section, we provide a concise review of the most pertinent studies in the field, encompassing datasets and methods for SGG in both natural imagery and SAI.
@@ 2.2 Scene Graph Generation Methods
In the following sections, we discuss existing methods for the SGG task and the challenges faced by the SGG task in large-size VHR SAI.
@@ 2.2.1 Scene Graph Generation Methods in Natural Imagery
Existing SGG methods for natural imagery have been dominated by the two-stage process consisting of OBD and relationship prediction between objects. Given the need for compatibility between OBD models and relationship prediction models within the SGG framework, the majority of existing SGG models, in alignment with the literature [43], adopt Faster R-CNN as the standard for the OBD task. To generate high-quality scene graphs from images, a series of works have been explored in several directions, such as task-specific [36, 44, 45, 46, 47] and dataset-specific [43, 48, 49]. Notably, aggregating contextual information between objects has been shown to be effective for SGG. Xu et al. [36] first combined contextual information to refine the features of objects and relationships by using GRU [50] to pass messages between graphs. Based on the discovery that there are some inherent relationship patterns in subgraphs, Motifs [44] constructed a new global context computing mechanism. A series of context-based studies [51, 52, 53, 54] have consecutively explored the messaging mechanisms of objects and relationships. Most recently, HetSGG [47] attempted to capture the relationships between objects in more detail by proposing a heterogeneous graph and proved its effectiveness. However, the above methods of exhaustive enumeration based on all pairs tend to result in excessive computation, which makes it difficult to run under common computational resources (e.g., 24 GB of memory in a single GPU) for large-size VHR SAI scenarios with a large number of objects and more object pairs.
@@ 4.1 Oriented Object Detection in Large-size VHR SAI
OBD in large-size SAI is an indispensable basis for SGG in large-size SAI, which directly affects the prediction accuracy of the triplets. Compared with small-size or low-resolution imagery, large-size VHR SAI can not only accommodate more complete large objects (e.g., runways, docks), but also greatly retain important details of small objects (such as the head and tail of a car, the head and tail of an airplane), which can help to mine more high-value knowledge in SAI. In processing large-size VHR SAI, mainstream deep learning-based OBD methods, whether supervised [56, 57, 58, 59] or weakly supervised [60, 61, 62, 63], adopt a cropping strategy [4, 64], which results in large objects, such as runways, being cut off. In addition to the cropping strategy, some OBD methods tackle the original large-size imagery with a fixed-window down-sampling strategy [65, 66, 67], resulting in the loss of massive detail information of the image. Recently, HBD-Net for single-class bridge detection in large-size VHR SAI is proposed [27], and the SDFF architecture in HBD-Net effectively solves the large scale variations in single-class bridge detection by fusing multi-scale contexts into the dynamic image pyramid (DIP) of the large-size image. However, this single-class OBD only considers the intra-class variations of the objects, which is not applicable in the multi-class OBD task with both intra-class and inter-class variations.
@@ 4.3.2 Prototype-guided Relationship Learning
LSKNet-S [97] CVPR’23 59.0 25.5 53.2 70.9 89.7 65.5 62.4 53.6 55.0 24.7 0.0 0.0 19.5 11.4 37.8
HOD-Net (Ours) - 56.1 19.5 48.4 64.2 89.6 63.9 50.5 57.4 55.7 20.5 12.6 30.2 23.0 66.8 45.6
@@
TABLE III: Results (%) of different training strategies on RSG test set. b_b denotes boarding_bridge, l_t denotes lattice_tower, s_l is ship_lock, and g_d represents gravity_dam. Swin-L [72] is adopted as the backbone.
@@ 5 Experiments and Analysis
In this subsection, we first introduce the sub-tasks related to the OBD and the SGG tasks, and then describe the evaluation metrics and our implementation details for the different tasks. Finally, extensive evaluation of our proposed methods for the OBD and SGG tasks are performed on the RSG dataset.
@@ 5.1 Benchmark
To evaluate the performance of the SGG model more comprehensively, the harmonized mean (HMR@K) of both MR@K and mMR@K is taken as the comprehensive evaluation metric for the SGG task in large-size VHR SAI. For all tasks, the prediction can be considered as a true positive (TP) result only when it matches the ground-truth triplet $<$subject, relationship, object$>$ and has at least 0.5 IoU between the bounding boxes of subject-object and the bounding boxes of ground-truth. As for the K, instead of the setting (K= 50/100), which is commonly used by the SGG task in natural imagery, we adopt the setting (K= 1000/1500). This is because, in large-size VHR SAI, the average number of triplets per image far exceeds that in natural imagery.
@@ 5.2 Implementation details
5.3 Results and Analysis for OBD
Comparison with Baselines. To explore the properties of RSG and provide guidelines for future OBD task in large-size VHR SAI, we conduct a comprehensive evaluation of about 30 OBD methods and analyze the results. Table II provides a comprehensive benchmark on RSG, including single/two-stage, CNN/Transformer, anchor-based/anchor-free and supervised/weakly-supervised methods. We adopt MMDetection [^1] https://github.com/Zhuzi24/RSG-MMDetection [100] and MMRotate [^2] https://github.com/yangxue0827/RSG-MMRotate [101] as the toolkits and all experiments are based on the standard ‘1x’ (12 epochs) training schedule. It can be observed that our HOD-Net significantly outperforms other methods under the same backbone (ResNet50 [73]) for both HBB-based and OBB-based detectors. To provide more accurate OBD results for the SGG task in large-size VHR SAI, we evaluated three OBD strategies on HBB-based and OBB-based detectors (see Table III): resizing, cropping, and our HOD-Net. It can be seen that our HOD-Net achieves 53.2% and 55.9% mAP on the HBB-based and OBB-based detectors, respectively (using a 0.5 IoU threshold). Specifically, under the same backbone (Swin-L [72]), it achieves a 10.6%/19.5% improvement in mAP compared to the baseline Faster R-CNN detector (cropping/resizing) in the HBB detection task, and a 12.9%/23.9% improvement in mAP metric compared to the baseline Oriented R-CNN detector (cropping/resizing) in the OBB detection task. Notably, for objects with extreme aspect ratios (e.g., runways, ship locks), our HOD-Net outperforms other methods in OBB detection by a great margin of (34.9%/31.2%). Above experiments indicate that our HOD-Net can provide more comprehensive and accurate object representation for SGG in large-size VHR SAI.
TABLE IV: Baseline results (%) of SGG on RSG test set. Our RPCM outperforms previous methods in all metrics on both HBB- and OBB-based detectors, especially on the more challenging SGCls and SGDet tasks. Specifically, our model outperforms previous state-of-the-art models by an absolute 5.17%/4.76% on HMR@1500/2000 for the SGCls task and 3.80%/3.80% on HMR@1500/2000 for the SGDet task on the OBB-based detector.
@@ 5.3 Results and Analysis for OBD
OBD
Type
Relationship
Prediction
Venue MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000
IMP [36] CVPR’17 49.90/51.62 18.51/19.40 27.00/28.20 42.48/43.63 16.46/17.02 23.73/24.49 14.98/15.03 5.11/5.13 7.62/7.65
Motif [44] CVPR’18 61.81/63.78 31.33/32.66 41.58/43.20 45.13/46.19 21.46/22.57 29.09/30.32 18.20/18.32 7.69/7.80 10.81/10.94
HETSGG [47] AAAI’23 63.83/65.26 28.70/29.72 39.60/40.84 30.03/31.08 10.50/11.05 15.56/16.30 18.11/18.18 5.91/5.99 8.91/9.01
PE-Net [46] CVPR’23 63.21/65.11 35.94/37.10 45.82/47.27 38.84/40.05 19.46/20.20 25.93/26.86 16.49/16.66 7.30/7.38 10.12/10.23
HBB RPCM (Ours) - 64.27/65.67 38.73/39.70 48.33/49.48 46.28/47.55 25.32/26.12 32.73/33.72 22.77/22.85 10.43/10.63 14.31/14.51
@@
TABLE V: Comparison of the PPG and other pair pruning methods on RSG test set. We report MR@1500/2000, mMR@1500/2000, and HMR@1500/2000(%) on PredCls, SGCls, and SGDet tasks.
@@
Qualitative Results and Visualization. In Fig. 8, we further give the visualizations of different strategies on HBB-based detectors and OBB-based detectors. It can be seen that many large objects (e.g., docks, aprons) cannot be detected holistically due to the integrity being destroyed in the detection based on cropping strategy, and a lot of small objects (e.g., airplanes, boats, cranes, boarding bridges) cannot be detected correctly due to the loss of information in the detection based on resizing strategy. The proposed HOD-Net shows superiority in the detection results of multi-scale objects and achieves almost the same results as the annotation for objects with extreme aspect ratios (e.g., runways). The above results show that the proposed HOD-Net can provide mandatory support for the SGG task in large-size VHR SAI.
@@ 5.4 Results and Analysis for SGG
Comparison with the State-of-the-art Methods. As shown in section 5.3, HOD-Net exhibits superior performance for the OBD task, which is a prerequisite for achieving accurate relationship prediction in the SGG task. For a fair comparison, all experiments utilized the same toolkit [^3] https://github.com/Zhuzi24/SGG-ToolKit and employed HOD-Net as the detector for the SGG task. We use MR@1500/2000, mMR@1500/2000, and HMR@1500/2000 as evaluation metrics of the SGG task. In Table IV, we conduct a comprehensive evaluation of 10 experiments and validated the effectiveness of RPCM by comparing it with four state-of-the-art SGG methods on both HBB-based and OBB-based detectors. As shown in Table IV, for the three sub-tasks of SGG, SGG based on the OBB-based detector shows better results overall, especially for the more challenging SGCls and SGDet tasks. For the SGCls task based on the OBB-based detector, our RPCM achieves 37.89%/38.92% on HMR@1500/2000, outperforming the SOTA method by a large margin of 5.17%/4.76%. For the SGDet task, our RPCM also outperforms the previous method with a 3.80%/3.80% improvement on HMR@1500/2000.
@@ 5.5 Ablation Study on SGG
In this section, we first explore the effect of PPG network, and then analyze the impact of PBA iters on SGG performance. Further ablation experiments are conducted on the RSG dataset to demonstrate the necessity of each key component in our proposed RPCM (i.e., object context augmentation (OCA), relationship context augmentation (RCA), and prototype matching).
Figure 11: Visualization of different PBA iters L. Blue edges represent the ground-truth relationships correctly predicted, purple edges indicate ground-truth relationships that failed to be detected, and green edges are reasonable relationships predicted by the model but not annotated in the ground-truth.
@@ 5.5.1 Effect of PPG Network
Table V gives the quantitative comparison results of the proposed PPG network with other pair pruning methods on the RSG dataset when the sorted top 10,000 pairs are selected. It can be seen that the ABBS method based on statistics has only a weak improvement over the randomized method in all metrics, which shows that the semantic and spatial combination types of the object pairs in large-size VHR SAI are so complex that it is difficult to obtain combination patterns of objects by statistics. The score matrix, as depicted in the first row of Fig. 10, becomes sparse after the PPG network learning, effectively retaining the annotated pairs. This highlights the robust learning capacity of the proposed PPG network when dealing with pairs containing rich knowledge. The local visualization shows that the PPG network can retain the annotated pairs with high scores, as well as some pairs that are non-annotated but may contain high-value relationships, this is because the annotation of triplets in the real annotation is hardly exhaustive, so achieving the same sparsity as ground-truth is not necessary.

### v2

@@ Abstract
Scene graph generation benchmark, large-size satellite imagery, object detection, relationship prediction.
@@ 2.2.1 Scene Graph Generation Methods in Natural Imagery
Existing SGG methods for natural imagery have been dominated by the two-stage process consisting of OBD and relationship prediction between objects. Given the need for compatibility between OBD models and relationship prediction models within the SGG framework, the majority of existing SGG models, in alignment with the literature [43, 44], adopt Faster R-CNN as the standard for the OBD task. To generate high-quality scene graphs from images, a series of works have been explored in several directions, such as task-specific [36, 45, 46, 47, 48] and dataset-specific [43, 49, 50]. Notably, aggregating contextual information between objects has been shown to be effective for SGG. Xu et al. [36] first combined contextual information to refine the features of objects and relationships by using GRU [51] to pass messages between graphs. Based on the discovery that there are some inherent relationship patterns in subgraphs, Motifs [45] constructed a new global context computing mechanism. A series of context-based studies [52, 53, 54] have consecutively explored the messaging mechanisms of objects and relationships. Most recently, HetSGG [48] attempted to capture the relationships between objects in more detail by proposing a heterogeneous graph and proved its effectiveness. However, the above methods of exhaustive enumeration based on all pairs tend to result in excessive computation, which makes it difficult to run under common computational resources (e.g., 24 GB of memory in a single GPU) for large-size VHR SAI scenarios with a large number of objects and more object pairs.
@@ 4.1 Oriented Object Detection in Large-size VHR SAI
OBD in large-size SAI is an indispensable basis for SGG in large-size SAI, which directly affects the prediction accuracy of the triplets. Compared with small-size or low-resolution imagery, large-size VHR SAI can not only accommodate more complete large objects (e.g., runways, docks), but also greatly retain important details of small objects (such as the head and tail of a car, the head and tail of an airplane), which can help to mine more high-value knowledge in SAI. In processing large-size VHR SAI, mainstream deep learning-based OBD methods, whether supervised [56, 57, 58, 59] or weakly supervised [60, 61, 62, 63], adopt a cropping strategy [4, 64], which results in large objects, such as runways, being cut off. In addition to the cropping strategy, some OBD methods tackle the original large-size imagery with a fixed-window down-sampling strategy [65, 66, 67], resulting in the loss of massive detail information of the image. Recently, HBD-Net for single-class bridge detection in large-size VHR SAI is proposed [27], and the SDFF architecture in HBD-Net effectively solves the large scale variations in single-class bridge detection by fusing multi-scale contexts via the dynamic image pyramid (DIP) of the large-size image. However, this single-class OBD only considers the intra-class variations of the objects, which is not applicable in the multi-class OBD task with both intra-class and inter-class variations.
@@ 4.3.2 Prototype-guided Relationship Learning
LSKNet-S [98] ICCV’23 59.0 25.5 53.2 70.9 89.7 65.5 62.4 53.6 55.0 24.7 0.0 0.0 19.5 11.4 37.8
PKINet-S [99] CVPR’24 55.2 22.4 48.0 67.5 89.9 59.1 56.3 50.7 47.4 21.1 0.0 0.0 19.0 0.2 32.8
HOD-Net (Ours) - 57.2 23.5 49.4 64.9 89.5 65.5 56.6 56.5 56.2 25.9 11.3 3.7 20.7 48.6 43.6
@@
TABLE III: Results (%) of different training strategies on RSG test set. b_b denotes boarding_bridge, l_t denotes lattice_tower, s_l is ship_lock, and g_d represents gravity_dam. All experiments are based on the standard ‘1x’ (12 epochs) training schedule. Swin-L [72] is adopted as the backbone.
@@ 5.1 Benchmark
To evaluate the performance of the SGG model more comprehensively, the harmonized mean (HMR@K) of both MR@K and mMR@K is taken as the comprehensive evaluation metric for the SGG task in large-size VHR SAI. For all tasks, the prediction can be considered as a true positive (TP) result only when it matches the ground-truth triplet $<$subject, relationship, object$>$ and has at least 0.5 IoU between the bounding boxes of subject-object and the bounding boxes of ground-truth. As for the K, instead of the setting (K= 50/100), which is commonly used by the SGG task in natural imagery, we adopt the setting (K= 1500/2000). This is because, in large-size VHR SAI, the average number of triplets per image far exceeds that in natural imagery.
@@ 5.2 Implementation details
TABLE IV: Baseline results (%) of SGG on RSG test set. Specifically, our model outperforms previous state-of-the-art models by an absolute 5.17%/4.76% on HMR@1500/2000 for the SGCls task and 1.27%/1.82% on HMR@1500/2000 for the SGDet task on the OBB-based detector.
@@
OBD Type Relationship Prediction Venue MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000
IMP [36] CVPR’17 49.90/51.62 18.51/19.40 27.00/28.20 42.48/43.63 16.46/17.02 23.73/24.49 23.01/23.72 8.25/8.60 12.15/12.62
Motif [45] CVPR’18 61.81/63.78 31.33/32.66 41.58/43.20 45.13/46.19 21.46/22.57 29.09/30.32 28.36/29.32 11.49/11.95 16.35/16.98
GPS-Net [53] CVPR’20 65.68/67.10 34.92/36.11 45.60/46.95 31.99/33.16 14.61/15.40 20.06/21.03 30.83/31.58 12.71/13.10 18.00/18.52
SHA [44] CVPR’22 64.03/65.77 31.98/33.16 42.66/44.10 48.08/49.54 21.60/22.46 29.81/30.91 30.21/31.04 13.24/13.65 18.41/18.96
HETSGG [48] AAAI’23 63.83/65.26 28.70/29.72 39.60/40.84 30.03/31.08 10.50/11.05 15.56/16.30 25.19/25.82 8.17/8.50 12.34/12.79
PE-Net [47] CVPR’23 63.21/65.11 35.94/37.10 45.82/47.27 38.84/40.05 19.46/20.20 25.93/26.86 28.11/29.22 12.80/13.40 17.59/18.37
HBB RPCM (Ours) - 64.27/65.67 38.73/39.70 48.33/49.48 46.28/47.55 25.32/26.12 32.73/33.72 30.33/31.36 14.09/14.78 19.24/20.09
@@
TABLE V: Comparison of our PPG and other pair pruning methods on RSG test set for the PredCls, SGCls, and SGDet tasks.
@@
5.3 Results and Analysis for OBD
Comparison with Baselines. To explore the properties of RSG and provide guidelines for future OBD task in large-size VHR SAI, we conduct a comprehensive evaluation of about 30 OBD methods and analyze the results. Table II provides a comprehensive benchmark on RSG, including single/two-stage, CNN/Transformer, anchor-based/anchor-free and supervised/weakly-supervised methods. We adopt MMDetection [^1] https://github.com/Zhuzi24/RSG-MMDetection [102] and MMRotate [^2] https://github.com/yangxue0827/RSG-MMRotate [103] as the toolkits and all experiments are based on the standard ‘1x’ (12 epochs) training schedule. It can be observed that our HOD-Net significantly outperforms other methods under the same backbone (ResNet50 [73]) for both HBB-based and OBB-based detectors. To provide more accurate OBD results for the SGG task in large-size VHR SAI, we evaluate three OBD strategies on HBB-based and OBB-based detectors (see Table III): resizing, cropping, and our HOD-Net. It can be seen that our HOD-Net achieves 53.2% and 55.9% mAP on the HBB-based and OBB-based detectors, respectively (using a 0.5 IoU threshold). Specifically, under the same backbone (Swin-L [72]), it achieves a 10.6%/19.5% improvement in mAP compared to the baseline Faster R-CNN detector (cropping/resizing) in the HBB detection task, and a 12.9%/23.9% improvement in mAP metric compared to the baseline Oriented R-CNN detector (cropping/resizing) in the OBB detection task. Notably, for objects with extreme aspect ratios (e.g., runways, ship locks), our HOD-Net outperforms other methods in OBB detection by a great margin of (34.9%/31.2%). Above experiments indicate that HOD-Net can provide more comprehensive and accurate object representation for SGG in large-size VHR SAI.
Qualitative Results and Visualization. In Fig. 8, we further give the visualizations of different strategies on HBB-based detectors and OBB-based detectors. It can be seen that many large objects (e.g., docks, aprons) cannot be detected holistically due to the integrity being destroyed in the detection based on cropping strategy, and a lot of small objects (e.g., airplanes, boats, cranes, boarding bridges) cannot be detected correctly due to the loss of information in the detection based on resizing strategy. The proposed HOD-Net shows superiority in the detection results of multi-scale objects and achieves almost the same results as the annotation for objects with extreme aspect ratios (e.g., runways). The above results show that HOD-Net can provide mandatory support for the SGG task in large-size VHR SAI.
@@ 5.4 Results and Analysis for SGG
Comparison with the State-of-the-art Methods. As shown in section 5.3, HOD-Net exhibits superior performance for the OBD task, which is a prerequisite for achieving accurate relationship prediction in the SGG task. For a fair comparison, all experiments utilized the same toolkit [^3] https://github.com/Zhuzi24/SGG-ToolKit and employed HOD-Net as the detector for the SGG task. We use MR@1500/2000, mMR@1500/2000, and HMR@1500/2000 as evaluation metrics of the SGG task. In Table IV, we conduct a comprehensive evaluation of 14 experiments and validated the effectiveness of RPCM by comparing it with six state-of-the-art SGG methods on both HBB-based and OBB-based detectors. As shown in Table IV, for the three sub-tasks of SGG, our RPCM shows better results overall, especially for the more challenging SGCls and SGDet tasks. For the SGCls task based on the OBB-based detector, our RPCM achieves 37.89%/38.92% on HMR@1500/2000, outperforming the SOTA method by a large margin of 5.17%/4.76%. For the SGDet task, our RPCM also outperforms the previous method with a 1.27%/1.82% improvement on HMR@1500/2000.
Figure 11: Visualization of different PBA iters L. Blue edges represent the correctly predicted relationships, purple edges indicate incorrectly predicted relationships, and green edges are reasonable relationships predicted by the model but not annotated in the ground-truth.
@@ 5.5.1 Effect of PPG Network
Table V gives the quantitative comparison of the proposed PPG network with other pair pruning methods on the RSG dataset when the sorted top 10,000 pairs are selected. It can be seen that the ABBS method based on statistics has only a weak improvement over the randomized method in all metrics, which shows that the semantic and spatial combination types of the object pairs in large-size VHR SAI are so complex that it is difficult to obtain combination patterns of objects by statistics. The score matrix, as depicted in the first row of Fig. 10, becomes sparse after the PPG network learning, effectively retaining the annotated pairs. This highlights the robust learning capacity of the proposed PPG network when dealing with pairs containing rich knowledge. The local visualization shows that the PPG network can retain the annotated pairs with high scores, as well as some pairs that are non-annotated but may contain high-value relationships, this is because the annotation of triplets in the real annotation is hardly exhaustive, so achieving the same sparsity as ground-truth is not necessary.